The most recent financial statements for Kerch, Inc., are shown here (assuming no income taxes): Income Statement Balance Sheet Sales $ 7,200 Assets $ 21,700 Debt $ 9,100 Costs 4,730 Equity 12,600 Net income $ 2,470 Total $ 21,700 Total $ 21,700 Assets and costs are proportional to sales. Debt and equity are not. No dividends are paid. Next year’s sales are projected to be $8,424. What is the external financing needed? (Do not round intermediate calculations and round your answer to 2 decimal places, e.g., 32.16.)
Answer #1

Growth rate in sales=(8424-7200)/7200=17%

Sales 8424
Less:costs(4730*1.17) $5534.10
Net income $2889.90

Total assets would be=$21700*117%=$25389

Total equity would be=$12600+$2889.90

=$15489.90

Total assets=Total equity +Total liabilities

Hence external financing needed=($25389-$15489.90-$9100)

which is equal to

=$799.10